Air is to be compressed in a diesel engine cylinder. The initial pressure and temperature are 105 kN/m2 and 15 °C respectively and the final pressure is 7 MPa. In order to ignite the fuel charge injected at the end of compression, the temperature required is 1100 °C. Calculate the ratio of the initial to the final volume for this to be possible.


1
Expert's answer
2021-05-30T13:28:44-0400

Initial Pressure, "P_1=105\times10^3\space Pa"

Final Pressure, "P_2=7\times10^{6}\space Pa"

Initial Temperature, "T_1=15\degree C=288\space K"

Final Temperature, "T_2=1100\degree C=1373\space K"

Initial Volume "=V_1"

Final Volume "=V_2"

"\dfrac{P_1V_1}{T_1}=\dfrac{P_2V_2}{T_2}"

"\dfrac{V_1}{V_2}=\dfrac{P_2\times T_1}{P_1\times T_2}"

"\dfrac{V_1}{V_2}=13.98"